1 out of 2 attemp A computer svstem administrator notices that computers runnine a particular operating svstem seem to freeze up more often as the installation of the operating svstem ages. She measures the time (in minutes) before freeze-up for 7 computers one month after installation, and for 9 computers seven months after installation. The results are as follows: One month after install: 207.2 233.3 215.2 234.9 224.6 243.9 245.2 Seven months after install: S3.S 52.S 126.3 200.9 173.7 245.8 149.3 155.S 102.4 Find a 95% confidence interval for the mean difference in time to freeze-up between the first month and the seventh Carry five sien ificant disits throuthout vour calculations
Answer #1

Solution:-

One month Seven month 229.1857 Mean 5.395211 Standard Error Mean Standard Error Median Mode Standard Deviation Sample Varianc

95% confidence interval for the mean difference in time to freeze up between the first month and seventh is C.I = (41.40909, 130.14091).

C.I = \left (\bar{x}_{1}-\bar{x}_{2} \right )\pm t_{\alpha /2}\times \sqrt{\frac{s_{1}^{2}}{n_{1}}+\frac{s_{2}^{2}}{n_{2}}}

C.I = (229.1857 - 143.1111) + 2.145 × 20.68341

C.I = 85.775 + 44.36591

C.I = (41.40909, 130.14091)